At Tuxi’s, tacos cost $2.50 (p=2.5) and the market wage for labor is $12.50. The optimal amount of labor is 6 workers.

Question: What would be the per-hour profits at Tuxi’s if capital costs $50 per hour?

Number of workers

Tacos

MPE

Average product

MRPE

Value of the AP

0

0

0

0

1

15

15

15

37.5

187.5

2

28

13

14

32.5

175

3

38

10

12.67

25.0

158.37

4

47

9

11.75

22.5

146.875

5

54

7

10.8

17.5

135.0

6

59

5

9.83

12.5

122.875

7

62

3

8.85

7.5

110.625

8

64

2

8.0

5.0

100.00

9

65

1

7.22

2.5

90.25
Answer #1

In the short run, the optimum point of production would be where the marginal revenue product of labor (MRPE) is equal to the price of labor ($12.5), which is occurring at L=6 and Q=59.

The variable cost would be L * 12.5= 6 * 12.5; dollar per hour, while the fixed cost is given as $50 per hour. The total cost would be hence the variable cost plus the fixed cost, ie $75+$50=$125.

The total revenue would be TR = PO or TR 2.5* 59= 147 dollars per hour.

The profit per hour would hence be T-TR _ TC 147.5-125 22 dollar per hour.